OpenStudy (anonymous):

simplify: 1+x+1-x+2sqrt(1-x^2)/(1-x)-(1-x)

OpenStudy (anonymous):

This has already been posted and answered for another student.

OpenStudy (anonymous):

2Sqrt(1-x^2) = 2Sqrt(1-x)(1+x) (1-x)+(1-x) = 2 -2x = 2(1-x) therefore now becomes... 2Sqrt(1-x)(1+x)/2(1-x) which simplifies to ... sqrt(1+x) now add the beginning part of the original expression insimlified form ...1+x+1-x =2 answer = 2+ sqrt(1+x)